# An Evaluation of Explanation Methods for Black-Box Detectors of Machine-Generated Text

### v1

@@ Abstract
The increasing difficulty to distinguish language-model-generated from human-written text has led to the development of detectors of machine-generated text (MGT). However, in many contexts, a black-box prediction is not sufficient, it is equally important to know on what grounds a detector made that prediction. Explanation methods that estimate feature importance promise to provide indications of which parts of an input are used by classifiers for prediction. However, the quality of different explanation methods has not previously been assessed for detectors of MGT. This study conducts the first systematic evaluation of explanation quality for this task. The dimensions of faithfulness and stability are assessed with five automated experiments, and usefulness is evaluated in a user study. We use a dataset of ChatGPT-generated and human-written documents, and pair predictions of three existing language-model-based detectors with the corresponding SHAP, LIME, and Anchor explanations. We find that SHAP performs best in terms of faithfulness, stability, and in helping users to predict the detector’s behavior. In contrast, LIME, perceived as most useful by users, scores the worst in terms of user performance at predicting the detectors’ behavior.
An Evaluation of Explanation Methods for Black-Box
Detectors of Machine-Generated Text
Loris Schoenegger           Yuxi Xia          Benjamin Roth Faculty of Computer Science, University of Vienna, Vienna, Austria
@@ 1 Introduction
Large language models, such as ChatGPT, produce output that is often virtually indistinguishable from human-written text. Their ability to generate human-like text at an unprecedented scale allows for new forms of phishing, disinformation campaigns, and academic fraud Crothers et al. (2023). Recent work has proposed language-model-based detection methods for machine-generated text (MGT, Solaiman et al., 2019; Guo et al., 2023; Mitchell et al., 2023). These operate as black-box detectors: they provide no explanations for their decisions. This is insufficient for applications that demand additional evidence, or when wrong decisions affect people, as would be the case for detecting MGT in academia.
@@
To address this, explanation methods have been applied to such detectors Mosca et al. (2023); Liu et al. (2023); Yu et al. (2023), with the majority of papers using SHAP Lundberg and Lee (2017) or LIME Ribeiro et al. (2016). We find that Anchor Ribeiro et al. (2018), a method that produces rule-based explanations, can also be applied in this setting. All three methods produce local explanations that explain single predictions (MGT or not) by locating relevant elements in the input (words in the document that influenced the prediction). Previously, LIME and SHAP have been used together with detectors of MGT without verifying the quality of the resulting explanations in this setting. However, if such explanations should aid in the categorization of a text as MGT or not, it must be verified that the explanation methods behave as intended. Moreover, reliable explanation methods can generate insights into differences that exist between human- and generated language, while unreliable explanation methods may generate a false sense of understanding of those differences.
In this work, we study explanation methods for detectors of MGT in three aspects of quality: Explanations should accurately depict the detector’s behavior (faithfulness: Jacovi and Goldberg, 2020; Ribeiro et al., 2016; Alvarez Melis and Jaakkola, 2018). They should be sensitive enough and sufficiently deterministic (stability: Alvarez Melis and Jaakkola, 2018; Lakkaraju et al., 2020; Nauta et al., 2023), and be effective at communicating the model’s decision process to users (usefulness: Hoffman et al., 2019; Doshi-Velez and Kim, 2017). To assess how suitable these methods ultimately are for explaining detector decisions, we perform a systematic evaluation with automated metrics and a user study. We test for the above-mentioned aspects with two existing fine-tuned Transformer-based detectors Guo et al. (2023); Solaiman et al. (2019) and a zero-shot method Mitchell et al. (2023).
@@
To enable a comparison of faithfulness, we adapt the ideas of a token removal experiment Arras et al. (2016) and the pointing game Poerner et al. (2018) to this task and all explanation methods, and we construct a dedicated test suite to measure the pointing game accuracy in this setting. Stability is assessed with controlled synthetic data checks Nauta et al. (2023) that we construct specifically for the task of detecting MGT. Previous user studies have evaluated usefulness on tasks for which humans already have an intuitive understanding, such as sentiment analysis or income prediction Ribeiro et al. (2016, 2018); Hase and Bansal (2020). Identifying patterns in a detector’s behavior is arguably more challenging for users when they have limited knowledge about the features that might be relevant. To keep forward simulation experiments feasible for humans in a setting where detectors may utilize complex and unintuitive feature sets, we use a study design similar to that of Hase and Bansal (2020) (originally from the context of movie review sentiment prediction and income prediction) but with a special document selection strategy. Rather than randomly selecting documents, we present users with pairs of documents that we choose based on explanation similarity.
Figure 1: Ranking in the experiments.
Figure 1 shows an overview of the results. In summary, the contributions of our work are: [^1] Code, data and explanations are made available at
https://github.com/loris3/evaluation_explanation_quality
  • (1)
@@
    The first study to systematically evaluate explanation methods for detectors of MGT, an important basis for understanding differences between generated and natural language; detectors should only be used in critical scenarios if accompanied by explanation methods for which strengths and limitations are experimentally validated.
  • (2)
@@
    Evaluation of all combinations of 3 detectors, 3 explanation methods with 5 automated experiments.
  • (3)
@@
  • (4)
    In addition to automated metrics, a user study to measure perceived usefulness and actual usefulness in a forward simulation.
  • (5)
    We find that SHAP performs best across all automated metrics. LIME explanations are perceived as most useful by the users but decrease users’ performance at predicting the decisions of the detectors. Neither Anchor nor LIME consistently outranks the other in our experiments.
@@ 2 Related Work
Evaluating Faithfulness Arras et al. (2016) evaluate faithfulness with token removal experiments. Note that this setup raises the same concern perturbation-based explanation methods do when applied to text classification: It assumes classifiers behave predictably for partial input.
Controlled synthetic data checks are an alternative type of experiment that foregoes this issue Nauta et al. (2023). Poerner et al. (2018) also proposed a similar design for evaluating local feature importance explanations for sentiment prediction. We extend both experiments to the task of explaining MGT detection, including rule-based Anchor explanations, in Section 3.1.
@@
Evaluating Stability Quantifying the stability of explanations across runs, consistency, with agreement measures is discussed in Nauta et al. (2023). The properties of continuity (similar explanations for similar documents with the same prediction, Alvarez Melis and Jaakkola, 2018) and contrastivity (sufficiently different explanations for similar documents but different predictions) described by Nauta et al. (2023) are alternative notions of stability that rely on convergence to a lesser extent. We propose a setup that generates coherent perturbations from a language model for these tests in Section 3.2.
@@
Evaluating Usefulness Hase and Bansal (2020) perform a forward simulation experiment Doshi-Velez and Kim (2017), where they assess whether users can predict the detector’s behavior better after being shown explanations. Hase and Bansal conduct their user study with two comparatively simple binary classifiers and tasks (sentiment analysis and income prediction). Given that the explanation methods analyzed here produce outcome- and not model explanations Guidotti et al. (2018), it is less likely that users will obtain a comprehensive understanding of the model’s behavior or relevant features and successfully apply that to new instances of human-written and MGT. We propose a strategy to increase the feasibility of this kind of experiment in Section 3.3. Besides measuring user performance, we also assess perceived usefulness. This is measured in a rating task with questions adapted from Hoffman et al. (2019).
3 Defining Explanation Quality Metrics for Detectors of MGT
@@
We systematically evaluate explanation quality along the axes of faithfulness, stability, and usefulness. Throughout this section, $f(d_{i})$ refers to the decision of a detector $f$ for a document $d_{i}$. $D$ is the base dataset of human-written and machine-generated documents $\{d_{i}\}$. SHAP and LIME provide feature importance scores for tokens from the input document $d_{i}$ as explanations for decisions made by $f$ Lundberg and Lee (2017). Those tokens that had the strongest impact on the detector’s decision should be attributed the highest scores. Anchor explanations are expressed as if-then rules Ribeiro et al. (2018). There can be multiple valid Anchors for a document. Ribeiro et al. (2018) define an Anchor $A_{i,m}$ for the document $d_{i}$ as a set of tokens $A_{i,m}=\{t_{i,j}\}$ that, if present in the document, guarantee the same decision as for $d_{i}$ with a probability greater than $\tau$ for perturbations in the local neighborhood of $d_{i}$.
@@ 3.1 Evaluation of Faithfulness
A faithful explanation method should accurately depict the detector’s behavior Jacovi and Goldberg (2020); Ribeiro et al. (2016); Alvarez Melis and Jaakkola (2018). The first test for this is a pointing game similar to Poerner et al. (2018). Random sentences from the base dataset $D$ are concatenated to form a synthetic dataset of hybrid documents $D^{h}=\{d_{i}^{h}\}$. The length of these hybrid documents is set to match the mean document length by the number of sentences in the original dataset in our setup. It is assumed that the detector’s decision $f(d_{i}^{h})$ on such a document is based on segments that were originally part of documents with ground truth $y_{i}$ equal to $f(d_{i}^{h})$. A faithful explanation method should hence find these segments to be more important for the decision than those originating from opposite-class documents. In the pointing game, feature importance explanation methods are therefore awarded hits for a document $d_{i}^{h}$ if the token with the highest feature importance score $t_{i,max}\in d_{i}^{h}$ was in fact originally part of a document with ground truth $y_{i}$ equal to $f(d_{i}^{h})$ Poerner et al. (2018): $S_{x}=\bigcup_{\forall d_{i}\in D}\{t\in d_{i}\,\mid\,y_{i}=x\}$, $hit(d_{i}^{h},t)=\mathbbm{1}[t\in S_{f(d_{i}^{h})}]$. Where $\mathbbm{1}[\cdot]$ is 1 if $\cdot$ evaluates to true, 0 otherwise.
The pointing game accuracy is the fraction of documents in the dataset $D^{h}$ that get awarded hits. For feature importance type explanations, it is given as: $\text{Acc}_{\text{pg}}=\sum\limits_{\forall d_{i}^{h}\in D^{h}}{hit(d_{i}^{h},t_{i,max})}/|D^{h}|$.
@@
Poerner et al. (2018) only evaluate feature-importance explanation methods. Note that there is no distinction by importance between tokens within an Anchor. Furthermore, a single Anchor can span multiple sentences. We therefore attribute hits proportionally for this explanation method. The hit function then tests all tokens $t_{i,j}$ specified by the Anchor $A_{i}$ individually and returns the average number of hits instead of a binary value: $hit^{R}(d_{i}^{h})=\frac{\sum_{\forall t_{i,j}\in A_{i}}hit(d_{i}^{h},t_{i,j})}{|A_{i}|}$.
@@
A token removal experiment as in Arras et al. (2016) is performed as the second test for faithfulness. Let $d_{i}$ be the original document and $d_{i}^{k}$ a version with the $k$ top-tokens by feature importance towards $f(d_{i})$ removed. Arras et al. plot the accuracy of the detector at different $k$ with respect to the ground truth $y_{i}$ of the original documents. For Anchor explanations, we remove tokens from the Anchor that applies to the highest proportion of documents in the local neighborhood (the one with the highest $\tau$) in random order.
@@ 3.2 Evaluation of Stability
We consider three different notions of stability. The individual experiments are based on the principles of consistency, continuity, and contrastivity as characterized in Nauta et al. (2023).
@@
Consistency We measure consistency across five explanations for the same document with the agreement metric Krippendorff’s $\alpha$ Krippendorff (1970), calculated on the explanations’ feature-importance vectors. [^2] Anchor explanations are one-hot encoded: If a token is part of an Anchor, it is deemed relevant.
Continuity The continuity between the explanation for the original document and explanations for a set of 5 perturbations $\{d_{i}^{\iota}\}$ is also measured with Krippendorff’s $\alpha$. For each document $d_{i}$ from the original dataset, a single token is randomly selected and replaced with an arbitrary number of tokens using the T5 language model (t5-small 60.5M params; Raffel et al., 2020). We set the maximum output length to 150 tokens. It is verified that $f(d_{i}^{\iota})=f(d_{i})$. In some instances, T5 fails to generate five unique replacements. In these cases, the token is replaced with a random token from the vocabulary.
@@
Contrastivity The high-level goal of contrastivity evaluation is to verify that documents that are similar in content, but are assigned different labels by the detector, get sufficiently different explanations Nauta et al. (2023). We implement this principle by editing documents from the dataset with a language model in order to obtain pairs of coherent documents $(d_{i},d_{i}^{\Omega})$ where $d_{i}$ and $d_{i}^{\Omega}$ are similar but get assigned different labels. In each iteration $k$, one token is deleted from the end of the document. The shortened document $d_{i}^{*}$ is used as an initial prefix for the autoregressive foundation model facebook/opt-350m Zhang et al. (2022) to generate a continuation that (together with the prefix) will constitute the perturbation $d_{i}^{\Omega}$. We halt removing tokens as soon as $f(d_{i}^{\Omega})\neq f(d_{i})$.
@@
To increase the chance of a label flip with smaller edits, five attempts are made at every $k$. Still, if the first tokens in the document are highly influential for the decision of the detector, more tokens need to be deleted until a label change can be observed, resulting in more dissimilar perturbed documents. [^3] E.g., documents that start with greetings like ”Hi!”. To retain a certain level of similarity between documents, perturbations that edit more than 50% of the original document are discarded. The resulting synthetic datasets of pairs (one set per detector) are further described in Figure 11 in Appendix D.
@@
Two scores are subsequently calculated on the explanations for $d_{i}$ and $d_{i}^{\Omega}$. These verify whether the explanations are consistent with the generation strategy: Given that the left part of $d_{i}$ and $d_{i}^{\Omega}$ are identical, but $f(d_{i}^{\Omega})\neq f(d_{i})$, one expects the filled-in part in $d_{i}^{\Omega}$ to have had a strong influence on the detector when labeling $d_{i}^{\Omega}$. A sufficiently sensitive explanation method should be able to detect this.
@@
The scores are formulated as synthetic data checks like the pointing game described above. $c_{\textit{inter}}$ is calculated on the parts that differ across the two documents, $d^{-}_{i}$ and $d^{\Omega-}_{i}$. It tests whether the mean feature importance score towards $f(d^{\Omega}_{i})$ (denoted $\mu_{\vec{v}^{~{}\Omega}}(\cdot)$) is higher for $d_{i}^{\Omega-}$ than it is for $d_{i}^{-}$. The hit function for this case is: $hit_{\textit{inter}}=\mathbbm{1}[\mu_{\vec{v}^{~{}\Omega}}(d_{i}^{\Omega-})>\mu_{\vec{v}^{~{}\Omega}}(d_{i}^{-})]$.
@@
$c_{\textit{intra}}$ shares the same intuition but is defined only on $d_{i}^{\Omega}$. As the left parts of $d_{i}$ and $d_{i}^{\Omega}$ are identical, but $f(d^{\Omega})\neq f(d_{i})$, one expects the generated part $d_{i}^{\Omega-}$ to have a higher average feature importance score towards $f(d_{i}^{\Omega})$ than the shared part: $hit_{\textit{intra}}=\mathbbm{1}[\mu_{\vec{v}^{~{}\Omega}}(d_{i}^{\Omega-})>\mu_{\vec{v}^{~{}\Omega}}(d_{i}^{\Omega*})]$.
@@
Both scores are again given as the fraction of documents in the dataset that score a hit: $c_{\{\textit{intra},\textit{inter}\}}=\sum\limits_{\forall(d_{i},d_{i}^{\Omega})}{hit_{\{\textit{intra},\textit{inter}\}}(d_{i},d_{i}^{\Omega})}/|D|$.
Figure 2: Information shown for LIME in phase 3 (left); and in the annotation phases 2 and 4 (right).

### v2

@@ Abstract
The increasing difficulty to distinguish language-model-generated from human-written text has led to the development of detectors of machine-generated text (MGT). However, in many contexts, a black-box prediction is not sufficient, it is equally important to know on what grounds a detector made that prediction. Explanation methods that estimate feature importance promise to provide indications of which parts of an input are used by classifiers for prediction. However, these are typically evaluated with simple classifiers and tasks that are intuitive to humans. To assess their suitability beyond these contexts, this study conducts the first systematic evaluation of explanation quality for detectors of MGT. The dimensions of faithfulness and stability are evaluated with five automated experiments, and usefulness is assessed in a user study. We use a dataset of ChatGPT-generated and human-written documents, and pair predictions of three existing language-model-based detectors with the corresponding SHAP, LIME, and Anchor explanations. We find that SHAP performs best in terms of faithfulness, stability, and in helping users to predict the detector’s behavior. In contrast, LIME, perceived as most useful by users, scores the worst in terms of user performance at predicting detector behavior.
An Evaluation of Explanation Methods for Black-Box Detectors of Machine-Generated Text
Loris Schoenegger1,2, Yuxi Xia1,2, Benjamin Roth1,3 1Faculty of Computer Science, University of Vienna, Vienna, Austria 2UniVie Doctoral School Computer Science, University of Vienna, Vienna, Austria 3Faculty of Philological and Cultural Studies, University of Vienna, Vienna, Austria Correspondence: loris.schoenegger@univie.ac.at
@@ 1 Introduction
Large language models, such as ChatGPT, produce output that is often virtually indistinguishable from human-written text. Their ability to generate human-like text at an unprecedented scale allows for new forms of phishing, disinformation campaigns, and academic fraud Crothers et al. (2023). Recent work has proposed language-model-based detection methods for machine-generated text (MGT; Solaiman et al., 2019; Guo et al., 2023; Mitchell et al., 2023). These operate as black-box detectors: they provide no explanations for their decisions. This is insufficient for applications that demand additional evidence, or when wrong decisions affect people, as would be the case for detecting MGT in academia.
@@
To address this, explanation methods have been applied to such detectors Mosca et al. (2023); Liu et al. (2023); Yu et al. (2023), with the majority of papers using SHAP Lundberg and Lee (2017) or LIME Ribeiro et al. (2016), two methods that assign feature importance scores to individual tokens in the input (higher importance scores suggest stronger impact on the detector’s decision). We find that Anchor Ribeiro et al. (2018), a method that generates rule-based explanations (if-then rules, e.g.,: If the document contains the word "delve", the detector will predict "machine" in 70% of cases), can also be applied in this setting. All three methods produce local explanations that explain single predictions (MGT or not) by locating relevant elements in the input (words in the document that influenced the prediction).
Previously, LIME and SHAP have been used together with detectors of MGT without verifying the quality of the resulting explanations. This is problematic as prior evaluations relied on tasks such as sentiment analysis, which have comparatively simple feature sets that are more intuitive to users. In this work, we rely on MGT detection as a case study for tasks characterized by understudied feature sets. The novelty of our work is an evaluation that goes beyond typical evaluation tasks, assessing explanations in terms of three important aspects of quality, defined as follows:
  1. (1)
    We investigate how accurately explanations can depict a detector’s behavior (faithfulness: Jacovi and Goldberg, 2020; Ribeiro et al., 2016; Alvarez Melis and Jaakkola, 2018). This is crucial as unfaithful explanations can create a false sense of understanding.
  2. (2)
    We study whether they are sensitive enough and sufficiently deterministic (stability: Alvarez Melis and Jaakkola, 2018; Lakkaraju et al., 2020; Nauta et al., 2023). Both aspects are prerequisites for building user trust.
  3. (3)
    And we investigate how effective they are at communicating the model’s decision process to users (usefulness: Hoffman et al., 2019; Doshi-Velez and Kim, 2017), as explanations must be able to generate meaningful insights.
Moreover, reliable explanation methods can generate insights into differences that exist between human- and generated language (e.g., syntactic structures, stylistic choices), while unreliable explanation methods may generate a false sense of understanding of those differences. To assess how suitable these methods ultimately are for explaining detector decisions, we perform a systematic evaluation with automated metrics and a user study. We test for the above-mentioned aspects with two existing fine-tuned Transformer-based detectors Guo et al. (2023); Solaiman et al. (2019) and a zero-shot method Mitchell et al. (2023).
@@
To enable a comparison of faithfulness, we adapt the ideas of a token removal experiment (where the effect that removing tokens selected by the explanation method has on the classifier’s decision is studied; Arras et al., 2016) and the pointing game (Section 3.1; Poerner et al., 2018) to this task and all explanation methods, and we construct a dedicated test suite to measure the pointing game accuracy in this setting. Stability is assessed with controlled synthetic data checks (which test if explanations are in line with a certain data generation strategy; Nauta et al., 2023) that we construct specifically for the task of detecting MGT. Note that previous user studies have evaluated usefulness on tasks for which humans already have an intuitive understanding, such as sentiment analysis or income prediction Ribeiro et al. (2016, 2018); Hase and Bansal (2020). Identifying patterns in a detector’s behavior is, however, arguably more challenging for users when they have limited knowledge about the features that might be relevant. To keep forward simulation experiments (in which users’ ability to anticipate the classifier’s behavior is measured before and after showing explanations) feasible for humans in a setting where detectors may utilize complex and unintuitive feature sets, we use a study design similar to that of Hase and Bansal (2020, originally from the context of movie review sentiment prediction and income prediction) but add a special document selection strategy: Rather than randomly selecting documents, we present users with pairs of documents that we choose based on explanation similarity.
Figure 1 presents the aggregated results of all our experiments. In summary, the contributions of our work are:
Figure 1: Scores in the experiments. Min-max normalized for each respective metric.
  1. (1)
@@
    The first study to systematically evaluate explanation methods for detectors of MGT, an important basis for understanding differences between generated and natural language; explanation methods can only be relied upon in critical scenarios if their strengths and limitations are experimentally validated.
  2. (2)
@@
    Evaluation of all combinations of three detectors, three explanation methods with five automated experiments.
  3. (3)
@@
  4. (4)
    In addition to automated metrics, a user study to measure perceived- and actual usefulness in a forward simulation experiment.
We find that SHAP performs best across all automated metrics. LIME explanations are perceived as most useful by the users but decrease users’ performance at predicting the decisions of the detectors. Neither Anchor nor LIME consistently outranks the other in our experiments. [^1] Code, data and explanations are made available at doi.org/10.5281/zenodo.15655989
@@ 2 Related work
Evaluating faithfulness.
Arras et al. (2016) evaluate faithfulness (the degree to which the explanation method accurately depicts the detector’s behavior; Jacovi and Goldberg, 2020; Ribeiro et al., 2016; Alvarez Melis and Jaakkola, 2018) with token removal experiments. This type of experiment tests whether removing tokens found to be important by the explanation method is more likely to change the classifier’s decision than removing random tokens. Note that this approach raises the same concern that perturbation-based explanation methods do when applied to text classification: It assumes that classifiers behave predictably for partial input, which is problematic for the task at hand as (accidental) omission of words might be an effective feature for distinguishing human-written from machine-generated text. Controlled synthetic data checks Nauta et al. (2023) are an alternative type of experiment that foregoes this issue. Poerner et al. (2018) propose a design of this type for evaluating local feature importance explanations for sentiment prediction. We extend both experiments to the task of MGT detection, including rule-based Anchor explanations, in Section 3.1.
Evaluating stability.
@@
Quantifying the stability of explanations across runs with agreement measures (consistency) is discussed in Nauta et al. (2023). The properties of continuity (similar explanations for similar documents with the same prediction; Alvarez Melis and Jaakkola, 2018) and contrastivity (sufficiently different explanations for similar documents but different predictions) described by Nauta et al. (2023) are alternative notions of stability that rely on convergence to a lesser extent. We propose a setup that generates coherent perturbations from a language model for these tests in Section 3.2.
Evaluating usefulness.
@@
Hase and Bansal (2020) perform a forward simulation experiment Doshi-Velez and Kim (2017), where they assess whether users can predict the detector’s behavior better after being shown explanations. Hase and Bansal conduct their user study with two comparatively simple binary classifiers and tasks (sentiment analysis and income prediction). Given that the explanation methods analyzed here produce outcome- and not model explanations (i.e., they explain one instance at a time; Guidotti et al., 2018), it is less likely that users will obtain a comprehensive understanding of the model’s behavior or relevant features and successfully apply that to new instances of human-written and MGT. We propose a strategy to increase the feasibility of this kind of experiment in Section 3.3. Besides measuring user performance, we also assess perceived usefulness. This is measured in a rating task with questions adapted from Hoffman et al. (2019).
3 Methods: Explanation Quality Metrics for MGT Detectors
@@
We systematically evaluate explanation quality along the axes of faithfulness, stability, and usefulness. Throughout this section, $f(d_{i})$ refers to the decision of a detector $f$ for a document $d_{i}$. $D$ is the base dataset of human-written and machine-generated documents $\{d_{i}\}$. SHAP and LIME provide feature importance scores for tokens from the input document $d_{i}$ as explanations for decisions made by $f$ Lundberg and Lee (2017). Those tokens that had the strongest impact on the detector’s decision should be attributed the highest scores. SHAP- and LIME explanations are largely similar from a user’s perspective, as they share the same explanation format and rely on similar visualization methods (e.g., feature importance plots: Figure 4 and Appendix A). We describe how these two methods differ in their calculation of feature importance in Section 4.3.
Anchor explanations are expressed as if-then rules. Ribeiro et al. (2018) define an Anchor $A_{i,m}$ for the document $d_{i}$ as a set of tokens $A_{i,m}=\{t_{i,j}\}$ that, if present in the document, guarantee the same decision as for $d_{i}$ with a probability greater than $\tau$ for perturbations in the local neighborhood of $d_{i}$. There can be multiple valid Anchors per document.
@@ 3.1 Evaluation of Faithfulness
A faithful explanation method should accurately depict the detector’s behavior Jacovi and Goldberg (2020); Ribeiro et al. (2016); Alvarez Melis and Jaakkola (2018). The first test for this is a pointing game similar to that in Poerner et al. (2018). We first describe the approach for feature importance explanations (SHAP and LIME) before extending this metric to rule-based explanations (Anchor).
Random sentences from the base dataset $D$ are concatenated to form a synthetic dataset of hybrid documents $D^{h}=\{d_{i}^{h}\}$ (Figure 2). In our setup, we set the number of sentences per hybrid document to the mean number of sentences per document in the original dataset, and sample random sentences without replacement until all of $D$ is used.
Figure 2: Hybrid documents in the pointing game. The detector’s prediction for a hybrid document $f(d^{h}_{i})$ is compared against the ground truth of $t_{i,max}$ (the token with the highest importance score in the explanation). Sentences labeled with m or h originate from machine-generated or human-written documents respectively
@@
It is assumed that the detector’s decision $f(d_{i}^{h})$ on a hybrid document is based on segments that were originally part of documents with ground truth $y_{i}$ equal to $f(d_{i}^{h})$. A faithful explanation method should hence find these segments to be more important for the decision than those originating from opposite-class documents. In the pointing game, feature importance explanation methods are therefore awarded hits for a document $d_{i}^{h}$ if the token with the highest feature importance score $t_{i,max}\in d_{i}^{h}$ was in fact originally part of a document with ground truth $y_{i}$ equal to $f(d_{i}^{h})$ Poerner et al. (2018):
$\begin{split}S_{x}=\bigcup_{\forall d_{i}\in D}\{t\in d_{i}\,\mid\,y_{i}=x\},\\ hit(d_{i}^{h},t)=\mathbf{1}[t\in S_{f(d_{i}^{h})}]\end{split}$ (1)
Where $x$ is a label (human or machine) predicted for a particular hybrid document so that $S_{human}$ is the set of all tokens from documents in $D$ with ground-truth humanhuman, and $\mathbf{1}[\cdot]$ is 1 if $\cdot$ evaluates to true, 0 otherwise. The pointing game accuracy is the fraction of documents in the dataset $D^{h}$ that get awarded hits. For feature importance type explanations, it is given as:
$\text{Acc}_{\text{pg}}=\sum\limits_{\forall d_{i}^{h}\in D^{h}}{hit(d_{i}^{h},t_{i,max})}/|D^{h}|$ (2)
@@
Poerner et al. (2018) only evaluate feature-importance explanation methods. We add support for rule-based explanations as follows: First, there is no distinction by importance between tokens within an Anchor. Furthermore, a single Anchor can span multiple sentences. We therefore attribute hits proportionally for this explanation method. The hit function then tests all tokens $t_{i,j}$ specified by the Anchor $A_{i}$ individually and returns the average number of hits instead of a binary value:
$hit^{R}(d_{i}^{h})=\frac{\sum_{\forall t_{i,j}\in A_{i}}hit(d_{i}^{h},t_{i,j})}{|A_{i}|}$ (3)
@@
We perform a token removal experiment as in Arras et al. (2016) as the second test for faithfulness. The assumption in this experiment is that the prediction is more likely to change if tokens important to the detector’s decision are removed, rather than random ones. For the two feature importance explanation methods (SHAP and LIME), let $d_{i}$ be the original document and $d_{i}^{k}$ a version with the $k$ top-tokens by feature importance towards $f(d_{i})$ removed. Arras et al. then plot the accuracy of the detector at different $k$ with respect to the ground truth $y_{i}$ of the original documents.
While rule-based Anchor explanations do not rank tokens by their importance, they still aim to select tokens that were highly important for the detector’s decision. To emulate the logic outlined above for this form of explanation, we remove tokens from the Anchor that applies to the highest proportion of documents in the local neighborhood (the one with the highest $\tau$) in random order. We perform token removal at increments of 1 for $k\in[0,10]$, at $k=15$, and then increments of 10 up to $k=100$.
@@ 3.2 Evaluation of Stability
As discussed in Section 2, stability has several definitions in the existing literature. We therefore propose metrics for three different notions of stability. Specifically, our experiments aim to assess the properties of consistency, continuity, and contrastivity as characterized in  Nauta et al. (2023). For simplicity, we describe these properties and our experiments for feature-importance explanation methods here. Anchor explanations are transformed into feature-importance explanations using one-hot encoding: if a token in a document is part of an Anchor, it is considered important.
Consistency.
@@
We measure consistency (stability of explanations between runs) across five explanations for the same document with the agreement metric Krippendorff’s $\alpha$ Krippendorff (1970), calculated on the explanations’ feature-importance vectors.
Continuity.
We measure continuity (similar explanations for similar documents with the same prediction; Alvarez Melis and Jaakkola, 2018) between the explanation for the original document and explanations for a set of 5 perturbations $\{d_{i}^{\iota}\}$ with Krippendorff’s $\alpha$ as well. For each $d_{i}$ from the original dataset, a single token is randomly selected and replaced with an arbitrary number of tokens using the T5 language model (t5-small 60.5M params: Raffel et al., 2020). It is verified that $f(d_{i}^{\iota})=f(d_{i})$ for all 5 perturbations $d_{i}^{\iota}$. We rely on T5 for generating synthetic documents because replacement with random tokens makes text appear more human-written to the detectors (likely due to decreased coherence). This makes the construction of valid perturbations through random replacement for machine-generated documents within a reasonable amount of attempts infeasible. While this adds machine-generated parts to all perturbations, we are only interested in and only calculate Krippendorff’s $\alpha$ on the matching parts of the original and perturbed documents. Note that we do not test for faithfulness in this experiment. This still aligns with the rationale behind the property of continuity, which states that the rest of the explanation should not drastically change if the classifier’s prediction remains the same (Nauta et al., 2023, regardless of the type of signal introduced). We set the maximum output length to 150 tokens. In some instances, T5 fails to generate five unique replacements. In these cases, the token is replaced with a random token from the vocabulary.
Contrastivity.
@@
The high-level goal of contrastivity evaluation is to verify that documents that are similar in content, but are assigned different labels by the detector, get sufficiently different explanations Nauta et al. (2023). We implement this principle by editing documents from the dataset with a language model in order to obtain pairs of coherent documents $(d_{i},d_{i}^{\Omega})$ where $d_{i}$ and $d_{i}^{\Omega}$ are similar but get assigned different labels. In each iteration $k$, one token is deleted from the end of the document (Figure 3). The shortened document $d_{i}^{*}$ is used as an initial prefix for the autoregressive foundation model facebook/opt-350m: Zhang et al. (2022) to generate a continuation that (together with the prefix) will constitute the perturbation $d_{i}^{\Omega}$. We halt removing tokens as soon as $f(d_{i}^{\Omega})\neq f(d_{i})$.
@@
To increase the chance of a label flip with smaller edits, five attempts are made at every $k$. Still, if the first tokens in the document are highly influential for the decision of the detector, more tokens need to be deleted until a label change can be observed, resulting in more dissimilar perturbed documents (e.g., documents that start with greetings like "Hi!"). To retain a certain level of similarity between documents, perturbations that edit more than 50% of the original document, are discarded and not used for calculating the metric. The resulting synthetic datasets of pairs (one set per detector) are further described in Appendix B.
@@
Two scores are subsequently calculated on the explanations for $d_{i}$ and $d_{i}^{\Omega}$. These verify whether the explanations are consistent with the generation strategy: Given that the left part of $d_{i}$ and $d_{i}^{\Omega}$ are identical, but $f(d_{i}^{\Omega})\neq f(d_{i})$, one expects the filled-in part in $d_{i}^{\Omega}$ to have had a strong influence on the detector when labeling $d_{i}^{\Omega}$. A sufficiently sensitive explanation method should detect this.
The scores are formulated as synthetic data checks, similar to the the pointing game described above. Note that unlike with that experiment, but similar to the experiment testing for continuity, it is sufficient here that the modification we apply to a given document causes the detector to predict the opposite label. This is because we only aim to verify whether the explanations exhibit a desired behavior (that tey appear sufficiently different if the documents are assigned different labels, not whether they are faithful), which does not require that the signal we introduce is of the opposite type.
@@
The first score, $c_{\textit{inter}}$ is calculated on the parts that differ across the two documents, $d^{-}_{i}$ and $d^{\Omega-}_{i}$. It tests whether the mean feature importance score towards $f(d^{\Omega}_{i})$ (denoted $\mu_{\vec{v}^{~\Omega}}(\cdot)$) is higher for $d_{i}^{\Omega-}$ than it is for $d_{i}^{-}$. The hit function for this case is:
$hit_{\textit{inter}}=\mathbf{1}[\mu_{\vec{v}^{~\Omega}}(d_{i}^{\Omega-})>\mu_{\vec{v}^{~\Omega}}(d_{i}^{-})]$ (4)
$c_{\textit{intra}}$
@@
shares the same intuition but is defined only on $d_{i}^{\Omega}$. As the left parts of $d_{i}$ and $d_{i}^{\Omega}$ are identical, but $f(d^{\Omega})\neq f(d_{i})$, one expects the generated part $d_{i}^{\Omega-}$ to have a higher average feature importance score towards $f(d_{i}^{\Omega})$ than the shared part:
$hit_{\textit{intra}}=\mathbf{1}[\mu_{\vec{v}^{~\Omega}}(d_{i}^{\Omega-})>\mu_{\vec{v}^{~\Omega}}(d_{i}^{\Omega*})]$ (5)
@@
Both scores are again given as the fraction of documents in the dataset that score a hit:
$c_{\{\textit{intra},\textit{inter}\}}=\sum\limits_{\forall(d_{i},d_{i}^{\Omega})}{hit_{\{\textit{intra},\textit{inter}\}}(d_{i},d_{i}^{\Omega})}/|D|$
Figure 3: Perturbation strategy for the contrastivity experiment.